OpenStudy (anonymous):

3. Find the derivative of ƒ(x) = x^2 + 5^x using the definition of a derivative (first principles).

OpenStudy (lgbasallote):

what's your answer here?

OpenStudy (lgbasallote):

here's a hint \[\frac{d}{dx} a^x = a^x \ln a\]

OpenStudy (anonymous):

I know the answer is 2x +5^x (log(5) through using the various derivative rules However, i dont know how I can show the log(5) part through first principles

OpenStudy (lgbasallote):

it's implicit differentiation

OpenStudy (lgbasallote):

have you discussed that lesson?

OpenStudy (anonymous):

No sir

OpenStudy (lgbasallote):

oh wait...definition as in limits?

OpenStudy (anonymous):

Yes first principles h->o (F(X+h)-f(x)/h

OpenStudy (lgbasallote):

uhmm maybe @zepp is more suitable for this job...

OpenStudy (lgbasallote):

sorry not good with limits..he is though...just wait for @zepp

OpenStudy (anonymous):

alright will do Im guessing by linking him he will be informed of this question existing? :p

OpenStudy (zepp):

Simply plug in the function and solve the limit? Where's the problem?

OpenStudy (anonymous):

How do you make it so that you get 5^x(log5)? I cant figure that out

OpenStudy (zepp):

Yep that's right

OpenStudy (zepp):

that's the derivative of 5^x

OpenStudy (zepp):

\[\large \frac{d}{dx}x^2=2x\] Cheating by using the fact that \(\huge \frac{d}{dx}x^n=nx^{(n-1)}\)

OpenStudy (anonymous):

Ive found it so that lim (h-0) is equal to 5^x(5^h-1)/h But how would I continue that? And I know thats the derivative, I just cant find the answer using first principles

OpenStudy (zepp):

So \[\large f(x)=x^2 + 5^x \\\large \frac{d}{dx}x^2 + 5^x=2x+5^x{\log(5)}\]

OpenStudy (anonymous):

5xlog(5) <------ How would I show that using first principles? For this specific question I cannot use any derivation rules

OpenStudy (zepp):

@Zarkon Any suggestion on deriving a function that consists of a constant raised to the x power?

OpenStudy (zepp):

Using the first principals?